Greenhab Report – February 20th

Crew 205 Green Hab Report – 20-FEB-2019

GreenHab Officer: Nathan Hadland

Environmental control: Heating and Cooling w/ ambient air (2.5 hrs)

Shade Cloth (40%)

Average temperatures:
Low: 17.3°C

High: 27.7°C

Hours of supplemental light: 5 Hours

Daily water usage for crops: 9 gallons

Water in Blue Tank: ~ 189 gallons

Crops watered at 0900 hours and 1800 hours:


1) 09.00 AM 19.8° C / 23 %
2) 12.00 PM 27.1° C / 20 %
3) 15:20 PM 19.3° C / 23 %
4) 18:00 PM 22.0° C / 30 %

Changes to crops: None.

Narrative: The nutrient supplementation in regolith experiments are still ongoing. However, due to mishaps in shipping, chlorophyll analysis may not be possible since the methanol has not arrived to the Hab (an organic solvent to help extract chlorophyll from the biomass). Consequently, these experiments may have to end without further analysis and they will be repeated upon return to Florida Institute of Technology using the same regolith samples. These experiments performed at the Hab do nevertheless represent the baseline tests astronauts will have to perform to determine the viability of substrates to support plant growth.

The nutrient deficiency observed in the tomato plants appear to be improving, since the yellow coloration is decreasing. The “Mystery Greens” from the spilled seed bank have sprouted and appear to be healthy, with two apparent variants of seedlings. Tomorrow, we will perform some additional grooming of dead biomass from several of the unhealthy plants to stimulate growth as well as contribute to our ISRU cellulosic ethanol study. Similar to yesterday, the door remained closed for most of the day due to the low ambient temperatures and overcast weather. However, at 1200 hours, the door was opened to allow for ambient airflow and closed at 1340 hours due to the drop in temperature.

Harvest: None.

Support/supplies needed: None.

Copyright © The Mars Society. All rights reserved. | Main Site